BRICS Research Series, 2006

June 1, 2007

This document is also available as PostScript and DVI.

References

[RS-06-19] Abstract, PostScript, PDF, DVI.

Michael David Pedersen. Logics for The Applied π Calculus. December 2006. viii+111 pp.

[RS-06-18] Małgorzata Biernacka and Olivier Danvy. A Syntactic Correspondence between Context-Sensitive Calculi and Abstract Machines. dec 2006. iii+39 pp. Extended version of an article to appear in TCS. Revised version of BRICS RS-05-22.

[RS-06-17] Abstract, PostScript, PDF, DVI.

Olivier Danvy and Kevin Millikin. A Rational Deconstruction of Landin’s J Operator. December 2006. ii+37 pp. Revised version of BRICS RS-06-4. A preliminary version appears in the proceedings of IFL 2005, LNCS 4015:55–73.

[RS-06-16] Abstract, PostScript, PDF.

Anders Møller. Static Analysis for Event-Based XML Processing. October 2006. 16 pp.

[RS-06-15] Abstract, PostScript, PDF, DVI.

Dariusz Biernacki, Olivier Danvy, and Kevin Millikin. A Dynamic Continuation-Passing Style for Dynamic Delimited Continuations. October 2006. ii+28 pp. Revised version of BRICS RS-05-16.

[RS-06-14] Abstract, PostScript, PDF, DVI.

Giorgio Delzanno, Javier Esparza, and Jiří Srba. Monotonic Set-Extended Prefix Rewriting and Verification of Recursive Ping-Pong Protocols. July 2006. 31 pp. To appear in ATVA ’06.

[RS-06-13] Abstract, PostScript, PDF, DVI.

Jiří Srba. Visibly Pushdown Automata: From Language Equivalence to Simulation and Bisimulation. July 2006. 21 pp. To appear in CSL ’06.

[RS-06-12] Abstract, PostScript, PDF, DVI.

Kristian Støvring. Higher-Order Beta Matching with Solutions in Long Beta-Eta Normal Form. June 2006. 13 pp. To appear in Nordic Journal of Computing, 2006.

[RS-06-11] Abstract, PostScript, PDF.

Kim G. Larsen, Ulrik Nyman, and Andrzej Wasowski. An Interface Theory for Input/Output Automata. June 2006. 40 pp. Appears in Misra, Nipkow and Sekerinski, editors, Formal Methods: 14th International Symposium, FM ’06 Proceedings, LNCS 4085, 2006, pages 82–97.

[RS-06-10] Abstract, PostScript, PDF.

Christian Kirkegaard and Anders Møller. Static Analysis for Java Servlets and JSP. June 2006. 23 pp. Full version of paper presented at SAS ’06.

[RS-06-9] Claus Brabrand, Robert Giegerich, and Anders Møller. Analyzing Ambiguity of Context-Free Grammars. April 2006. 19 pp.

[RS-06-8] Abstract, PostScript, PDF.

Christian Kirkegaard and Anders Møller. Static Analysis for Java Servlets and JSP. April 2006. 22 pp.

[RS-06-7] Abstract, PostScript, PDF, DVI.

Petr Jančar and Jiří Srba. Undecidability Results for Bisimilarity on Prefix Rewrite Systems. April 2006. 20 pp. Presented at FoSSaCS 2006, LNCS 3921:277–291.

[RS-06-6] Abstract, PostScript, PDF, DVI.

Luca Aceto, Willem Jan Fokkink, Anna Ingólfsdóttir, and Bas Luttik. A Finite Equational Base for CCS with Left Merge and Communication Merge. March 2006. 22 pp.

[RS-06-5] Abstract, PostScript, PDF, DVI.

Kristian Støvring. Extending the Extensional Lambda Calculus with Surjective Pairing is Conservative. March 2006. 18 pp. To appear in Logical Methods in Computer Science. Supersedes RS-05-35.

[RS-06-4] Abstract, PostScript, PDF, DVI.

Olivier Danvy and Kevin Millikin. A Rational Deconstruction of Landin’s J Operator. February 2006. ii+26 pp. To appear in the post-reviewed proceedings of the 17th International Workshop on the Implementation and Application of Functional Languages (IFL’05), Dublin, Ireland, September 2005.

[RS-06-3] Abstract, PostScript, PDF, DVI.

Małgorzata Biernacka and Olivier Danvy. A Concrete Framework for Environment Machines. February 2006. ii+29 pp. To appear in the ACM Transactions on Computational Logic. Supersedes BRICS RS-05-15.

[RS-06-2] Abstract, PDF.

Mikkel Baun Kjærgaard and Jonathan Bunde-Pedersen. A Formal Model for Context-Awareness. February 2006. 26 pp.

[RS-06-1] Abstract, PostScript, PDF, DVI.

Luca Aceto, Taolue Chen, Willem Jan Fokkink, and Anna Ingólfsdóttir. On the Axiomatizability of Priority. January 2006. 25 pp.